The Creighton family owns a vacation rental which generates $10,809 in income each quarter. The family outsources maintenance to another company which charges $1,127 per quarter, with the amount increasing by $25 each subsequent quarter. What is the equivalent uniform quarterly worth of the vacation rental when looked at over a 15-year planning period? Use a nominal annual interest rate of 2%, compounded quarterly to make the calculation.

Enter your answer as 12345

Round your answer. Do not use a dollar sign ("$"), any commas (","), or a decimal point (".").
Answer #1

Ans: 8982

Explanation:

Here,

n = 4 * 15 years = 60

i = 2% / 4 = 0.5%

Equivalent uniform quarterly worth

= 10,809 - [1127 + 25(A/G, 0.5%, 60)]

= 10,809 - [1127 + 25(28.006)]

= 10,809 - 1,827.15

= 8982